OpenStudy (anonymous):

Which ordered pair is a solution of the inequality? y=> 4x-5 A. (3,4) B.(2,1) C.(3,0) D.(1,1)

OpenStudy (phi):

each pair is (x,y) to test if a pair "works" replace x and y in the equation with the numbers in the pair for example for A (3,4) x is 3 and y is 4. use these numbers in y ≥ 4*x-5 4 ≥ 4*3 -5 now do the arithmetic and see if the statement is true. in this case, the right side is 4*3-5 = 12-5= 7 4 ≥ 7 is NOT true. So option A is out. can you test the other choices ?

OpenStudy (anonymous):

Got it D. Thank you